Session: Feb 2001 (session 1)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Tonino Pisanu at Mon Feb 5 12:00:31 MET 2001) Success

Torun (K.Borkowski at Mon Feb 5 12:12:59 MET 2001) Minor failures
Only second half (since 10:32) of experiment recorded because of the recorder failure just after initial setup of the system. Luckily the failure has been traced to a bad (burned) connector in the electronics of tape drive. Uncooled receivers were used for this test.

Jodrell_Bank (Alastair Gunn at Mon Feb 5 12:52:16 MET 2001) Success
No known problems. Fringes between Mk2 - Cm seen in MERLIN.

Cambridge (Alastair Gunn at Mon Feb 5 12:52:35 MET 2001) Success
No known problems.

Noto (C.Trigilio at Tue Feb 6 13:57:12 MET 2001) Minor failures
Some tracking problem may occur due to a fault in one of the two elevation motors. No other problem reported.

Effelsberg (A. Kraus at Thu Feb 8 09:44:12 MET 2001) Failure
We did not participate!

Onsala (John Conway at Mon Feb 12 10:01:02 MET 2001) Failure
Did not participate - telescope mechanical problem.